what is the product of this (x+3)(3x-5)?
Answer by smik(40)

To solve the question, we have to multiply (x+3) by (3x-5). To do this we do something called factoring. In this case, we would multiply x by 3x, and x by -5. Then we would multiply 3 by 3x, and 3 by -5. So, let's do it:


x*3x = 3x^2
x*(-5) = -5x
3*(3x) = 9x
3*(-5) = -15


Therefore, the final answer would like this (after adding the like terms: -5x+9x):


3x^2 + 4x - 15